# On the non-hermitian Kitaev chain

### v1

@@ Abstract
We study the non-hermitian Kitaev chain model, for arbitrary complex parameters. In particular, we give a concise characterisation of the curves of eigenvalues in the complex plane in the infinite size limit. Using this solution, we characterise under which conditions the skin effect is absent, and for which eigenstates this is the case. We also fully determine the region in parameter space for which the model has a zero mode.
@@ I Introduction
The study of non-hermitian systems, in various contexts, has been extremely intense during the last years. These studies cover the properties of non-hermitian systems in general (focussing on the differences and similarities with hermitian systems), studies of particular non-hermitian models, as well as utilising non-hermitian systems for actual applications, using various different types of physical systems BeBuKu2019 ; BrLiBeRoAr2022unpub ; Br2014 ; DeYoHa2021 ; Ed2022 ; EdKuBe2019 ; GoAsKaTaHiUe2018 ; KaBeSa2019 ; KaShUeSa2019 ; KoBu2020 ; KuDw2019 ; KuEdBuBe2018 ; MaBe2021 ; MaVaTo2018 ; LaQv2023 ; Le2016 ; LeLiGo2019 ; LeBlHuChNo2017 ; LiWaDuCh2024unpub ; LiWeWuRuChLeNi2024unpub ; LiLi2024 ; LiTaYaLiLe2023 ; LiLiAn2023 ; MaAuPePaAc2024 ; MoArBe2023 ; NeRo2024 ; OkKaShSa2020 ; OrHe2023 ; RaSiRaJi2024unpub ; RiRo2022 ; RoOrAgHe2024 ; SaCiCABa2024 ; ShZhFu2018 ; StBe2021 ; VyRo2021 ; YaBe2024unpub ; YaMa2023 ; YaMoBe2023 ; YaWa2018 ; YoPeKaHa2019 ; XiDeWaZhWaYiXu2019 ; Xi2018 ; ZeBe2024 ; ZhSoZhZhLiCaLi2024 ; ZhLe2019 . One aspect that is of particular relevance for the current paper is the presence of the skin-effect, which is closely related to the breakdown of the famed bulk boundary correspondence of hermitian topological systems HeBaRe2019 ; Lo2019 ; LeTh2019 ; LiZhAiGoKaUeNo2019 ; FlZoVaBeBaTi2022 ; ZiReRo2021 ; BoKrSl2020 ; YoMu2019 ; YaMoBe2021 ; EdKuYoBe2020 ; Sc2020 ; BrHy2019 ; EdAr2022 .
In this paper, we focus solely on Kitaev chain Ki2001 famous for its Majorana zero modes in its topological phase (see Al2012 for a review). We consider the non-hermitian version of this model, which has, in various incarnations, been studied before Yu2017 ; KlCaDaMaWu2017 ; LiZhZhSo2018 ; LiJiSo2020 ; KaAsKaUe2018 ; Li2019 ; YaSo2020 ; ZhGuKoZhLi2021 ; SaNiTa2022 ; SaLa2023 ; ShSo2023 ; LiWaSoWa2024 ; RaRoKuKaSa2023 ; Zh2024 ; MoNa2024 ; FuKaMo2024 ; PiMeAg2024unpub ; CaAg2024unpub ; LiLiZhTiArLi2024 . We concisely characterise the location of the eigenvalues in the complex plane in the limit for an open chain in the infinite system size. It is clearly natural to ask for which general choice of parameters does the non-hermitian Kitaev chain have a zero-mode. Interestingly, the full answer to this question is arguably a bit more complicated than one might expect.
@@
In the current paper, we provide explicit answers to the questions we posed above for the non-hermitian Kitaev chain. We introduce the non-hermitian Kitaev chain, in order to set the notation, and start with the general analysis to solve the model in Sec. II. We continue in Sec. III by providing the solution of the non-hermitian Kitaev chain, for chains of finite length, under the restriction that the left and right hopping parameters $t_{1}$ and $t_{2}$ are equal. In Sec. IV we give a complete characterisation of the parameters, for which the (non-hermitian) model has a zero-mode. Here, we directly work in the infinite system size limit. In Sec. V, we provide a concise characterisation of the eigenvalues of the model, again for fully generic parameters, in the infinite system size limit. Using this characterisation, one can easily obtain the curves in the complex plain corresponding to the eigenvalues of the model. In Sec. VI, we analyse the solution to determine under which conditions the model exhibits skin effect, and if this is the case, to which eigenvalues this pertains. In Sec. VII we briefly discuss our results.
@@ II The non-hermitian Kitaev chain
In this section we state the model we are interested in, namely the Kitaev chain with nearest neighbour hopping and pairing, for arbitrary, complex parameters and start with the analysis to find the eigenvalues. We are mainly interested in open chains, with ‘free’ boundary conditions. In this paper, we do not study the transition from the open to the periodic system. In terms of fermion creation and annihilation operators, $c_{j}^{\dagger}$, $c_{j}$, the model for $L$ sites reads
@@
This ansatz is inspired by the ansatz used in the hermitian case, where $a$ in general is a phase (due to the superconducting nature of the model). Because we deal with the non-hermitian case, $a$ will have arbitrary modulus. The ‘bulk’ equations now take the following form,
@@ III Solution for finite $L$, with $t_{2}=t_{1}$ or $d_{1}d_{2}=0$
Before continuing the analysis of the model for infinite system size, we briefly consider the case with either $d_{1}=0$ or $d_{2}=0$, but not necessarily both, and otherwise arbitrary parameters (so we allow $t_{1}\neq t_{2}$ here). In this case, the model is closely related to the Hatano-Nelson model HaNe1996 ; HaNe1997 ; HaNe1998 , for which an exact solution that interpolates between open and periodic boundary conditions was presented in EdAr2022 . Using the techniques of that paper, we find that the eigenvalues when $d_{2}=0$, the eigenvalues do not depend on $d_{1}$ (and the other way around), but a subset of the eigenvectors does depend on $d_{1}$. The eigenvalues are given by
@@
In this section, we determine under which conditions the model has a zero mode. We consider arbitrary complex parameters, and work directly in the $L\rightarrow\infty$ limit. In finite systems, the energy of a zero mode decays exponentially with system size, but because we work in the limit $L\rightarrow\infty$, we put $\lambda=0$ identically. For $\lambda=0$, it turns out that one can obtain the solutions of the bulk equations, Eq. (7), explicitly
@@ V Solution in the $L\rightarrow\infty$ limit
This observation allows us to obtain a rather compact representation of the eigenvalues $\lambda$ in the limit of large system size, $L\rightarrow\infty$. Namely, we demand that two of the roots $x_{i}$ of the bulk equation Eq. (8) have the same absolute value. That is, we write
@@
The condition $x_{1}x_{2}x_{3}x_{4}=1$ is one of the Vieta equations, relating the roots of a polynomial to its coefficients Vi1646 . The remaning three Vieta equations for the roots can then be written as
@@
The curve(s) in the complex plane determined by the eigenvalues $\lambda$ in the limit $L\rightarrow\infty$ are obtained by varying $0\leq\alpha<2\pi$ (we note that $\alpha$ plays the role of the momentum $k$ in the periodic case). In principle, one can eliminate $\kappa$ and $s$ from the Vieta equations, to obtain an equation for $\lambda$ in terms of $\alpha$ and the parameters of the model. This results in a forth order equation in $\lambda^{2}$, which is not insightful. In practise, if one wants to obtain the actual curve, one can simply solve the Eqs. (31) numerically. It should be noted that not all solutions for $\lambda$ correspond to actual eigenvalues. This is because it needs to be checked that the roots are ordered as $|x_{1}|\geq|x_{2}|=|x_{3}|\geq|x_{4}|$. There are two branches for which $|x_{2}|=|x_{3}|\geq|x_{1}|,|x_{4}|$ or $|x_{1}|,|x_{4}|\geq|x_{2}|=|x_{3}|$, which do not lead to actual eigenvalues. Thus, we need the solutions of the Vieta Eqs. (31) such that $1/|s|\leq|\kappa|^{2}\leq|s|$. Equivalently, one can instead require that $|s|\leq|\kappa|^{2}\leq 1/|s|$.
Figure 1: Plot for the non-hermitian Kitaev Chain with parameters $m=3/2$, $t_{1}=i$, $t_{2}=2$, $d_{1}=d_{2}=3$. The black lines correspond to the eigenvalues of the infinite system. The blue and yellow lines correspond to solutions for $\lambda$ of the Vieta equations, that do not correspond to actual eigenvalues of the model. The green dots correspond to eigenvalues of the finite system with $L=100$.
@@ VI Presence of the skin effect
Because we have a rather compact characterisation of the eigenvalues of the model in the thermodynamic limit we can, in principle, analyse in detail for which parameters the model has a skin effect. In solving the model in the infinite size limit, we obtained that two roots of the bulk equation are equal in absolute value, see Eq. (30). Because the roots $x_{2}$ and $x_{3}$ with $|x_{2}|=|x_{3}|$ determine the corresponding eigenvector, we find that there is no skin effect when $|x_{2}|=|x_{3}|=1$. The eigenvalues of the eigenstates that do not show a skin effect, lie on the curves of the eigenvalues in the periodic case $\lambda_{\pm}(k)$, as given by Eq. (11). We are interested in determining the parameters of the model, for which there are eigenstates that do not show the skin effect for extended ranges of $k$. We will not in general try to locate isolated points.
There is a long history of determining the location of the roots of polynomials in the complex plane and several algorithms exist to determine, say, how many roots have a negative real part, without having to determine the actual roots. Such algorithms are often used in stability analyses of various systems. For instance, the Routh–Hurwitz stability criterion Ro1877 ; Hu1895 determines if all the roots of a polynomial have negative real parts. One way to derive the criterion is to construct the sequence of Sturm polynomials associated with the polynomial under investigation St1829 .
@@
Figure 3: The amplitude of eigenstate coefficients for two states of the finite $L=100$ model with parameters $m=2/5$, $t_{1}=2$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$. The left panel corresponds to a real eigenvalue ($\lambda\approx 3.0182$) showing skin effect (note the logarithmic scale), the right panel corresponds to a purely imaginary eigenvalue ($\lambda\approx 4.3949i$) without skin effect.
Figure 4: The eigenvalues for two sets of parameters. The colour coding is the same as in Fig. 2. The parameters for the left panel are $m=0$, $t_{1}=2$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$; in this case, none of the eigenstates exhibit skin effect, thus the red and gray curves coincide. The purely imaginary eigenvalues extend to ±$\pm(1+2\sqrt{3})i\approx\pm 4.4641i$. We cut off the figure for clarity. The parameters for the right panel are $m=1/2$, $t_{1}=2$, $t_{2}=1$, $d_{1}=d_{2}=\sqrt{3}$; in this case, all the eigenstate exhibit skin effect. We note that in this case, there is zero-mode.
In Fig. 2, we plot the eigenvalues of the system with parameters $m=2/5$, $t_{1}=2$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$. The figure clearly shows that for these parameters (which do no fall in one of the classes $m=0$ or $t_{1}=\pm t_{2}$), some of the eigenstates of the infinite open chain do show skin effect, while others do not. The eigenvalues that are purely imaginary do not show skin effect, while those that lie in the region bounded by the two ovals (corresponding to eigenvalues of the periodic case) do show skin effect. Interestingly, there are real (non-zero) eigenvalues that correspond to states that do exhibit skin effect. In the hermitian case, non-zero (and necessarily real) eigenvalues do not exhibit skin effect. In Fig. 3, we plot the absolute value of the eigenstate coefficients as a function of position for two eigenvalues of the finite chain namely $\lambda\approx 3.0182$ in the left panel (real eigenvalue with skin effect, using a logarithmic scale) and $\lambda\approx 4.3949i$ in the right panel (purely imaginary eigenvalue without skin effect).

### v2

@@ Abstract
We study the non-hermitian Kitaev chain model, for arbitrary complex parameters. In particular, we give a concise characterisation of the curves of eigenvalues in the complex plane in the infinite size limit, using a novel method which can be applied to other non-hermitian systems. Using this solution, we characterise under which conditions the skin effect is absent, and for which eigenstates this is the case. We also fully determine the region in parameter space for which the model has a zero mode.
@@ I Introduction
The study of non-hermitian systems, in various contexts, has been extremely intense during the last years. These studies cover the properties of non-hermitian systems in general (focussing on the differences and similarities with hermitian systems), studies of particular non-hermitian models, as well as utilising non-hermitian systems for actual applications, using various different types of physical systems AgKoLaSe2024 ; AgKoLaSe2023unpub ; BeBuKu2019 ; BrLiBeRoAr2022unpub ; Br2014 ; DeYoHa2021 ; Ed2022 ; EdKuBe2019 ; Ez2019 ; Ez2024 ; GoAsKaTaHiUe2018 ; KaBeSa2019 ; KaShUeSa2019 ; KoBu2020 ; KuDw2019 ; KuEdBuBe2018 ; Ma2024a ; Ma2024b ; MaBe2021 ; MaRo2023 ; MaVaTo2018 ; LaQv2023 ; Le2016 ; LeLiGo2019 ; LeBlHuChNo2017 ; LiWaDuCh2024unpub ; LiWeWuRuChLeNi2024unpub ; LiLi2024 ; LiTaYaLiLe2023 ; LiLiAn2023 ; MaAuPePaAc2024 ; MoArBe2023 ; NeRo2024 ; OkKaShSa2020 ; OrHe2023 ; RaSiRaJi2024unpub ; RiRo2022 ; RoOrAgHe2024 ; SaCiCABa2024 ; ShZhFu2018 ; StBe2021 ; VyRo2021 ; YaBe2024unpub ; YaMa2023 ; YaMoBe2023 ; YaWa2018 ; YoPeKaHa2019 ; XiDeWaZhWaYiXu2019 ; Xi2018 ; ZeBe2024 ; ZhSoZhZhLiCaLi2024 ; ZhLe2019 . One aspect that is of particular relevance for the current paper is the presence of the skin-effect, which is closely related to the breakdown of the famed bulk boundary correspondence of hermitian topological systems HeBaRe2019 ; Lo2019 ; LeTh2019 ; LiZhAiGoKaUeNo2019 ; FlZoVaBeBaTi2022 ; ZiReRo2021 ; BoKrSl2020 ; YoMu2019 ; YaMoBe2021 ; EdKuYoBe2020 ; Sc2020 ; BrHy2019 ; EdAr2022 .
In this paper, we focus solely on Kitaev chain Ki2001 famous for its Majorana zero modes in its topological phase (see Al2012 for a review). We consider the non-hermitian version of this model, which has, in various incarnations, been studied before Yu2017 ; KlCaDaMaWu2017 ; LiZhZhSo2018 ; LiJiSo2020 ; KaAsKaUe2018 ; Li2019 ; YaSo2020 ; ZhGuKoZhLi2021 ; SaNiTa2022 ; SaLa2023 ; ShSo2023 ; LiWaSoWa2024 ; RaRoKuKaSa2023 ; Zh2024 ; MoNa2024 ; FuKaMo2024 ; PiMeAg2024unpub ; CaAg2024unpub ; LiLiZhTiArLi2024 . A main result is that we concisely characterise the location of the eigenvalues in the complex plane in the limit for an open chain in the infinite system size. Using our novel method, we can explain why the curves the eigenvalues lie on exhibit branch points. Using our solution, we can address several other interesting questions.
It is clearly natural to ask for which general choice of parameters does the non-hermitian Kitaev chain have a zero-mode. Interestingly, the full answer to this question is arguably a bit more complicated than one might expect.
@@
In the current paper, we use our novel method to obtain the eigenvalues in the thermodynamic limit in order to provide explicit answers to the questions we posed above for the non-hermitian Kitaev chain. We introduce the non-hermitian Kitaev chain, in order to set the notation, and start with the general analysis to solve the model in Sec. II. We continue in Sec. III by providing the solution of the non-hermitian Kitaev chain, for chains of finite length, under the restriction that the left and right hopping parameters $t_{1}$ and $t_{2}$ are equal. In Sec. IV we give a complete characterisation of the parameters, for which the (non-hermitian) model has a zero-mode. Here, we directly work in the infinite system size limit. In Sec. V, we provide a concise characterisation of the eigenvalues of the model, again for fully generic parameters, in the infinite system size limit. Using this characterisation, one can easily obtain the curves in the complex plain corresponding to the eigenvalues of the model. In Sec. VI, we analyse the solution to determine under which conditions the model exhibits skin effect, and if this is the case, to which eigenvalues this pertains. In Sec. VII we summarise our results by means of an example, followed by a short discussion in Sec. VIII.
@@ II The non-hermitian Kitaev chain
In this section we state the model we are interested in, namely the Kitaev chain Ki2001 with nearest neighbour hopping and pairing, for arbitrary, complex parameters and start with the analysis to find the eigenvalues. We are mainly interested in open chains, with ‘free’ boundary conditions. In this paper, we do not study the transition from the open to the periodic system. In terms of fermion creation and annihilation operators, $c_{j}^{\dagger}$, $c_{j}$, the model for $L$ sites reads
@@
If the model were hermitian, $m$ would take the role of the chemical potential, $t_{1}$ and $t_{2}$ would be the hopping parameters, while $d_{1}$ and $d_{2}$ would correspond to the superconducting order parameter.
@@
This ansatz is inspired by the ansatz used in the hermitian case, where $a$ in general is a phase (due to the superconducting nature of the model). Because we deal with the non-hermitian case, $a$ will have arbitrary modulus. The structure of the powers of $x$ is the standard ansatz used in solving the type of recurrence relations for the bulk of systems with periodicity. In the physics literature, this dates back at least the seminal paper of Lieb, Schultz and Mattis LiScMa1961 . Using this ansatz, the ‘bulk’ equations take the following form,
@@
These eigenvalues corresponding to the model with periodic boundary conditions are obtained by first performing a (discrete) Fourier transform. This results in the following $2\times 2$ matrix
$\begin{pmatrix}m+t_{1}e^{ik}+t_{2}e^{-ik}&d_{1}(e^{ik}+e^{-ik})\\ -d_{2}(e^{ik}+e^{-ik})&-m-t_{1}e^{-ik}-t_{2}e^{ik}\end{pmatrix}\ .$ (12)
The eigenvalues of this matrix are indeed given by Eq. (11).
@@ III Solution for finite $L$, with $t_{2}=t_{1}$ or $d_{1}d_{2}=0$
Before continuing the analysis of the model for infinite system size, we briefly consider the case with either $d_{1}=0$ or $d_{2}=0$, but not necessarily both, and otherwise arbitrary parameters (so we allow $t_{1}\neq t_{2}$ here). In this case, the model is closely related to the Hatano-Nelson model HaNe1996 ; HaNe1997 ; HaNe1998 , for which an exact solution that interpolates between open and periodic boundary conditions was presented in EdAr2022 . Using the techniques of that paper, we find that for $d_{2}=0$, the eigenvalues do not depend on $d_{1}$ (and the other way around), but a subset of the eigenvectors does depend on $d_{1}$. The eigenvalues are given by
@@
In this section, we determine under which conditions the model has a zero mode. Just as in the hermitian Kitaev chain Ki2001 , the zero modes exhibit an exponential decay away from the boundaries, and correspond to topological edge states.
We consider arbitrary complex parameters, and work directly in the $L\rightarrow\infty$ limit. In finite systems, the energy of a zero mode decays exponentially with system size, but because we work in the limit $L\rightarrow\infty$, we put $\lambda=0$ identically. For $\lambda=0$, it turns out that one can obtain the solutions of the bulk equations, Eq. (7), explicitly
@@ V Solution in the $L\rightarrow\infty$ limit
This observation allows us to obtain a rather compact representation of the eigenvalues $\lambda$ in the limit of large system size, $L\rightarrow\infty$. Namely, we demand that two of the roots $x_{i}$ of the bulk equation Eq. (8) have the same absolute value.
That is, we write
@@
The condition $x_{1}x_{2}x_{3}x_{4}=1$ is one of the Vieta equations, relating the roots of a polynomial to its coefficients Vi1646 . Before continuing our analysis, we quickly explain the Vieta equations. For ease of presentation, we do this for a fourth order polynomial, as the generalisation to the arbitrary case is clear. We write the polynomial in terms of its coefficients $a_{i}$ and its roots $r_{j}$ as follows
$P(x)=a_{4}x^{4}+a_{3}x^{3}+a_{2}x^{2}+a_{1}x+a_{0}=a_{4}\prod_{j=1}^{4}(x-r_{j})\ .$ (32)
Expanding the right hand side, and comparing with the coefficients leads to the following equations
$r_{1}+r_{2}+r_{3}+r_{4}$ $=-a_{3}/a_{4}$ (33)
$r_{1}r_{2}+r_{1}r_{3}+r_{1}r_{4}+r_{2}r_{3}+r_{2}r_{4}+r_{3}r_{4}$ $=a_{2}/a_{4}$ (34)
$r_{1}r_{2}r_{3}+r_{1}r_{2}r_{4}+r_{1}r_{3}r_{4}+r_{2}r_{3}r_{4}$ $=-a_{1}/a_{4}$ (35)
$r_{1}r_{2}r_{3}r_{4}$ $=a_{0}/a_{4}\ .$ (36)
These are the Vieta equations, which relate the coefficients of a polynomial to its roots. We note that the left hand sides correspond to the elementary symmetric polynomials evaluated at the roots of the original polynomial $P(x)$.
In the case at hand we have $a_{0}=a_{4}$, leading to the fourth Vieta equation $x_{1}x_{2}x_{3}x_{4}=1$ as stated above. The remaning three Vieta equations for the roots can then be written (after taking a linear combination) as
@@
The curve(s) in the complex plane determined by the eigenvalues $\lambda$ in the limit $L\rightarrow\infty$ are obtained by varying $0\leq\alpha<2\pi$ (we note that $\alpha$ plays the role of the momentum $k$ in the periodic case). In principle, one can eliminate $\kappa$ and $s$ from the Vieta equations, to obtain an equation for $\lambda$ in terms of $\alpha$ and the parameters of the model. This results in a forth order equation in $\lambda^{2}$, which is not insightful. In practise, if one wants to obtain the actual curve, one can simply solve the Eqs. (37) numerically. It should be noted that not all solutions for $\lambda$ correspond to actual eigenvalues. This is because it needs to be checked that the roots are ordered as $|x_{1}|\geq|x_{2}|=|x_{3}|\geq|x_{4}|$. There are two branches for which $|x_{2}|=|x_{3}|\geq|x_{1}|,|x_{4}|$ or $|x_{1}|,|x_{4}|\geq|x_{2}|=|x_{3}|$, which do not lead to actual eigenvalues. Thus, we need the solutions of the Vieta Eqs. (37) such that $1/|s|\leq|\kappa|^{2}\leq|s|$. Equivalently, one can instead require that $|s|\leq|\kappa|^{2}\leq 1/|s|$. Despite the fact that there are two non-physical branches, that do not lead to actual eigenvalues of the model, these non-physical branches play an important role in explaining the geometry of the actual curves the eigenvalues lie on. Without the additional constraint that the roots need to satisfy $|x_{1}|\geq|x_{2}|=|x_{3}|\geq|x_{4}|$, one would obtain smooth curves for the eigenvalues in the complex plane. With the additional necessary constraint however, one finds that the actual curves on which the eigenvalues lie exhibit branching points. This is because these actual curves ‘jump from one branch to another’.
Figure 1: Plot for the non-hermitian Kitaev Chain with parameters $m=3/2$, $t_{1}=i$, $t_{2}=2$, $d_{1}=d_{2}=3$. The black lines correspond to the eigenvalues of the infinite system. The blue and yellow lines correspond to solutions for $\lambda$ of the Vieta equations, that do not correspond to actual eigenvalues of the model (as explained in the main text). The green dots correspond to eigenvalues of the finite system with $L=100$.
@@ VI Presence of the skin effect
Because we have a rather compact characterisation of the eigenvalues of the model in the thermodynamic limit we can, in principle, analyse in detail for which parameters the model has a skin effect. Eigenvectors that exhibit skin effect are exponentially localised to the boundary of the systems. We determine the parameters for which the system exhibits a skin effect, and if so, to which eigenvectors this pertains (there can be cases where only a subset of eigenvectors exhibit skin effect).
In solving the model in the infinite size limit, we obtained that two roots of the bulk equation are equal in absolute value, see Eq. (31). Because the roots $x_{2}$ and $x_{3}$ with $|x_{2}|=|x_{3}|$ determine the corresponding eigenvector, we find that there is no skin effect when $|x_{2}|=|x_{3}|=1$. The eigenvalues of the eigenstates that do not show a skin effect, lie on the curves of the eigenvalues in the periodic case $\lambda_{\pm}(k)$, as given by Eq. (11). We are interested in determining the parameters of the model, for which there are eigenstates that do not show the skin effect for extended ranges of $k$. We will not in general try to locate isolated points.
There is a long history of determining the location of the roots of polynomials in the complex plane. One of the main reasons for this, is that it is used extensively in system analysis, in particular for systems that are linear and time invariant. Such a system is stable, if the output is bounded, even in the limit when the evolved time goes to infinity. This is the case when all the roots of the characteristic polynomial corresponding to the system have negative real parts. Several algorithms exist to determine, how many roots have a negative real part, without having to determine the actual roots. Routh and Hurwitz independently developed such algorithms, leading to the Routh–Hurwitz stability criterion Ro1877 ; Hu1895 , which determines if all the roots of a polynomial have negative real parts. One way to derive the criterion is to construct the sequence of Sturm polynomials associated with the polynomial under investigation St1829 .
@@
We do not repeat the full Bistritz algorithm here (which can be found in Bi2002 ), because it is a bit lengthy, and we do not need most of the details for our purposes. Therefor, we focus instead on the parts of the algorithm that are relevant for our analysis.
@@
Figure 3: The amplitude of eigenstate coefficients $|\psi_{i}|^{2}$ as a function of the site index $i$ for two states of the finite $L=100$ model with parameters $m=2/5$, $t_{1}=2$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$. The blue squares (orange dots) correspond to the first (second) component of the eigenvector of a given site $i$. The left panel (a) corresponds to a real eigenvalue ($\lambda\approx 3.0182$) showing skin effect (note the logarithmic scale), the right panel (b) corresponds to a purely imaginary eigenvalue ($\lambda\approx 4.3949i$) without skin effect.
Figure 4: The eigenvalues for two sets of parameters. The colour coding is the same as in Fig. 2. The parameters for the left panel (a) are $m=0$, $t_{1}=2$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$; in this case, none of the eigenstates exhibit skin effect, thus the red and gray curves coincide. The purely imaginary eigenvalues extend to ±$\pm(1+2\sqrt{3})i\approx\pm 4.4641i$. We cut off the figure for clarity. The parameters for the right panel (b) are $m=1/2$, $t_{1}=2$, $t_{2}=1$, $d_{1}=d_{2}=\sqrt{3}$; in this case, all the eigenstate exhibit skin effect. We note that in this case, there is zero-mode.
In Fig. 2, we plot the eigenvalues of the system with parameters $m=2/5$, $t_{1}=2$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$. The figure clearly shows that for these parameters (which do no fall in one of the classes $m=0$ or $t_{1}=\pm t_{2}$), some of the eigenstates of the infinite open chain do show skin effect, while others do not. The eigenvalues that are purely imaginary do not show skin effect, while those that lie in the region bounded by the two ovals (corresponding to eigenvalues of the periodic case) do show skin effect. Interestingly, there are real (non-zero) eigenvalues that correspond to states that do exhibit skin effect. In the hermitian case, non-zero (and necessarily real) eigenvalues do not exhibit skin effect. In Fig. 3, we plot the absolute value of the eigenstate coefficients as a function of position for two eigenvalues of the finite chain namely $\lambda\approx 3.0182$ in the left panel (a) (real eigenvalue with skin effect, using a logarithmic scale) and $\lambda\approx 4.3949i$ in the right panel (b) (purely imaginary eigenvalue without skin effect). In particular, in the left panel of Fig. 3, we find that the structure of the eigenvector is an exponentially damped oscillation. The exponential decay signifies the presence of the skin effect. In terms of the non-Bloch band theory of non-hermitian systems introduced in YoMu2019 , the eigenvectors that do not exhibit skin effect have a $\beta$ parameter such that $|\beta|=1$ (or a real momentum), while the eigenvectors that do exhibit skin effect have a $\beta$ parameter such that $|\beta|\neq 1$ (i.e., its ‘momentum’ is complex).